OpenStudy (anonymous):

Hey Can Someone Either please Message me or Comment If Your Really Good With Algebra, i need some help, i think i got it, but will someone go over it with me? If so Thanks <3(i'll fan)^.^

OpenStudy (mathmale):

Courtney, I'd suggest you post your question(s) in the usual way.

OpenStudy (anonymous):

Please Do Act Like i Havent Tried (:

OpenStudy (anonymous):

I'm sorry but I gotta but my input... @mathmale You shouldn't be like that... people only want medals and fans so you have to do it that way you don't because your a moderator and of you have any problems you can feel free to message me and @Courtneydyrdek98 just ignore him

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

and for a tip... none of the moderators listen to you and your welcome I feel the same way you do when moderators think their running something you have

OpenStudy (mathmale):

In the time it took you to type these unhelpful comments, we could have gotten started on the problem you wanted to solve, and perhaps solved it. Sorry you feel that way about moderators. What evidence have you that they don't listen to you?

OpenStudy (anonymous):

Goodbye @mathmale

OpenStudy (anonymous):

Part A: Solve A = 9 over 2(x + 28) for x. (4 points) Part B: Determine the value of x when A = 135. (2 points) Part C: Solve –np – 40 > 10 for n. Show your work. (4 points)

OpenStudy (anonymous):

YOU I was once a fan of you until you did this to some one else and I disliked that so I unfanned you... and I'm pretty sure she can look the answer up somewhere else so were done here @mathmale you can leave now and @Courtneydyrdek98 you should close the question and look up the answer on a different website

OpenStudy (anonymous):

I'm sorry you had a bad experience with mathmale so did I

OpenStudy (anonymous):

ive tried to /:

OpenStudy (anonymous):

@AccessDenied can you help us @312856MLP @bakonloverk @hoblos

OpenStudy (anonymous):

and really mathmale we don't care what you have too say

OpenStudy (anonymous):

it's just and un helpful comment

OpenStudy (anonymous):

i just need help :'( </3

OpenStudy (mathmale):

\[\frac{ A }{ 1 }=\frac{ 9 }{2(x+28) }\] can be inverted to make it easier to solve for x:\[\frac{ 1 }{ A }=\frac{ 2(x+28) }{ 9 }\] Multiply both sides by 9: \[\frac{ 9 }{ A }=2(x+28)\]Divide both sides by 2: \[\frac{ 9 }{ 2A }=x+28\] Can you now solve for x?

OpenStudy (anonymous):

i dont understand any of that.. /.\

OpenStudy (mathmale):

Do you want to understand it? If so, ask for clarification of whatever it is that you don't understand. Unfortunately there is no simpler way to solve this problem.

OpenStudy (anonymous):

its fine...

OpenStudy (mathmale):

Sorry your search for help isn't working out. I'd suggest you repost your very first question, separately, and delete this conversation. A fresh post, especially a simple one, is more likely to attract responses. Good luck.

OpenStudy (anonymous):

thank you , im sorry about earlier and everything.

OpenStudy (mathmale):

And I also. Hope next time we meet on OpenStudy that all goes well; I'll do my best to see that it does.

OpenStudy (anonymous):

i know i just rejected your help , but i need it now /;

OpenStudy (mathmale):

How can I help in a way that is least painful for you? Care to go over the explanation I typed in earlier? Thank you for re-considering.

OpenStudy (anonymous):

i understand it i think now that i read it , from 9/2a = x+28 how would i find x?

OpenStudy (mathmale):

Just subtract 28 from both sides. What do you now get for x?

OpenStudy (anonymous):

how do i subtract 28 from 9/2a?

OpenStudy (mathmale):

Just write it as\[\frac{ 9 }{ 2A }-28\]

OpenStudy (mathmale):

Not worth trying to simplify that.

OpenStudy (anonymous):

so what would x be?

OpenStudy (mathmale):

\[x=\frac{ 9 }{2A}-28\]

OpenStudy (anonymous):

okay so now , part b?

OpenStudy (mathmale):

Let me look at it.

OpenStudy (anonymous):

okay thank you

OpenStudy (mathmale):

We are to determine the value of x when A=135. Just go back to the most recent equation. Write in ' 135 ' in place of A. Simplify the fraction. Write out your result: \[x=\frac{ 9 }{ 2*135 }-28\]Simplify the result as much as possible (but don't sweat it).

OpenStudy (anonymous):

i dont know how i would simplify this /.\

OpenStudy (mathmale):

\[\frac{ 9 }{ 2*135 }=\frac{ 9 }{ 270 }=\frac{ 9 }{9*30 }\]

OpenStudy (mathmale):

What say you cancel out the 9? What's left?

OpenStudy (anonymous):

what do you mean?

OpenStudy (mathmale):